# Enhancing Adversarial Robustness of Vision Language Models via Adversarial Mixture Prompt Tuning

### v1

@@ Abstract
Large pre-trained Vision Language Models (VLMs) have excellent generalization capabilities but are highly susceptible to adversarial examples, presenting potential security risks. To improve the robustness of VLMs against adversarial examples, adversarial prompt tuning methods are proposed to align the text feature with the adversarial image feature without changing model parameters. However, when facing various adversarial attacks, a single learnable text prompt has insufficient generalization to align well with all adversarial image features, which finally leads to the overfitting phenomenon. To address the above challenge, in this paper, we empirically find that increasing the number of learned prompts can bring more robustness improvement than a longer prompt. Then we propose an adversarial tuning method named Adversarial Mixture Prompt Tuning (AMPT) to enhance the generalization towards various adversarial attacks for VLMs. AMPT aims to learn mixture text prompts to obtain more robust text features. To further enhance the adaptability, we propose a conditional weight router based on the input adversarial image to predict the mixture weights of multiple learned prompts, which helps obtain sample-specific aggregated text features aligning with different adversarial image features. A series of experiments show that our method can achieve better adversarial robustness than state-of-the-art methods on 11 datasets under different experimental settings.
@@ 1 Introduction
Large pre-trained Vision Language Models (VLMs) such as CLIP [27] have excellent generalization capabilities and can be regarded as foundation models [2] in different downstream tasks, e.g., image-text retrieval, zero-shot image classification, or image generation guidance. Due to its wide range of application scenarios, it places high requirements on security performance. However, despite its excellent performance, VLMs face many potential security risks [13, 24, 28], including the fact that visual models are vulnerable to adversarial examples [31], which can generate serious threat to the application in actual scenarios.
@@
To eliminate this potential security risk, many works have been proposed to improve the robustness of VLMs to adversarial example, which can be mainly divided into two types, full-parameter fine-tuning [24, 32, 29, 35] and parameter-efficient fine-tuning [24, 32, 36, 17, 39, 21]. Among them, full-parameter fine-tuning is an effective method to improve the adversarial robustness of the model. However, this method often requires a lot of computational overhead and also affects the performance of the model on general tasks. Another type of parameter-efficient method, e.g., adversarial prompt tuning [36], freezes all or most of the weights of the model and only fine-tunes some of its parameters. This type of method can also effectively improve the adversarial robustness with lower training overhead compared with full-parameter fine-tuning, which is an ideal solution. However, adversarial prompt tuning may face a serious problem: insufficient generalization. For example, for the text prompt tuning [36, 17], when only one learnable prompt is fine-tuned, the text feature is not sufficient to fit the image features for various adversarial examples, which can easily lead to overfitting and further cause potential security risks [33].
@@
Figure 1: The framework of Adversarial Mixture Prompt Tuning (AMPT). To enhance the adversarial robustness, we apply adversarial mixture prompt to generate diverse individual text feature, and utilize the conditional prompt weight router to obtain a sample-specific mixture text feature, and finally bring more generalization towards different adversarial examples.
@@
To enhance the generalization of the adversarial text prompt, an intuitive approach is to increase the length of the text prompt. However, we find that when it grows to a certain length, a longer prompt will bring greater optimization difficulty, and also needs higher requirements on the corresponding text encoder to deal with long prompts, finally leading to suboptimal robustness. Inspired by the Mixture of Experts (MoE) paradigm [4], we adopt an alternative strategy: increasing the number of learnable base prompts. Similar to multiple experts in MoE, we construct a composite prompt by combining several base prompts, with their weights adaptively determined based on the characteristics of the adversarial image. This approach enables the generation of more diverse and expressive text features. Moreover, since each base prompt remains short and is easier to optimize, leading to improved adversarial robustness. A preliminary experiment has empirically verified our idea, that is, increasing prompt number can indeed bring stronger robustness than increasing prompt length.
@@
Based on the above consideration, in this paper, we propose an adversarial prompt tuning method named Adversarial Mixture Prompt Tuning (AMPT) to enhance the adversarial robustness of VLMs. Specifically, we fix the parameters of the text and image encoders but only optimize adversarial mixture prompts. These text prompts pass through the text encoder and generate diverse individual text features. In addition, in order to enhance the adaptability, we propose a conditional text weight router based on image features to predict the weights of adversarial mixture prompts and aggregate them into a sample-specific aggregated text feature, so as to adaptively align with the diverse adversarial image features. A series of experiments show that our AMPT can achieve better accuracy and robustness than state-of-the-art methods on multiple different datasets. Meanwhile, AMPT also shows better generalization across different datasets. Our contribution can be summarized as follows:
@@
  • We find that for adversarial text prompt tuning, increasing the number of learnable text prompts can achieve a better robustness improvement than only increasing the length of learnable text prompts within a certain range of parameters.
@@
  • We propose a novel method named Adversarial Mixture Prompt Tuning (AMPT), which applies adversarial mixture prompts to generate diverse individual text features, where each text feature can play its unique roles for different adversarial examples, thereby alleviating the overfitting phenomenon.
@@
  • We apply a conditional text weight router based on image features to predict the weights of different text features and obtain a sample-specific aggregated text feature that has pretty adaptability to align with different adversarial image features. Furthermore, we theoretically verify the effectiveness of the weight router.
@@
  • We empirically verify the effectiveness of AMPT. Extensive experiments demonstrate that our AMPT can outperform state-of-the-art methods against adversarial examples in adversarial robustness and generalization across different datasets.
@@ 2 Related Work
2.1 Prompt Tuning for Accuracy in Vision-Language Models
@@
2.2 Adversarial Prompt Tuning in Vision-Language Models
Due to its excellent performance and low training cost, prompt tuning has been applied to improve the robustness of VLMs. [5] applies visual prompting to enhance the adversarial robustness. Furthermore, TeCoA [24] and PMG-AFT [32] employ visual prompt tuning to improve the adversarial robustness of VLMs. AdvPT [36] and APT [17] are proposed to apply text prompt tuning to further enhance the VLMs against image attacks. FAP [39] tries to enhance the robustness via bimodal tuning, while APD [21] further extends FAP into the adversarial distillation setting. To solve the insufficient generalization, [33] applies Test-Time Adversarial Prompt Tuning (TAPT) to learns defensive bimodal (textual and visual) prompts. Different from the above research, in this paper, we try to improve the adversarial robustness via adversarial mixture prompt tuning in the training process.
3 The Necessity of Adversarial Mixture Prompts
3.1 Formulation of Adversarial Prompt Tuning in CLIP
@@
where $x$ and $t$ are the image and text pairs belong to the dataset $\mathcal{D}$. For the image classification task with $N$ classes, texts $t$ also contain $N$ different prompts: $\{t_{1},t_{2},\cdots,t_{N}\}$. $\tilde{x}$ denotes adversarial examples. $y$ denotes the ground truth. $y_{ij}$ indicates whether the image $x_{i}$ and text $t_{j}$ pair match, if the image $x_{i}$ and text $t_{j}$ match, $y_{ij}$ is equal to 1, otherwise $y_{ij}$ is equal to 0; the $F_{\theta_{v}}$ and $F_{\theta_{t}}$ are the image encoder and text encoder of CLIP.
Meanwhile, as for the text $t$, a fixed text template, e.g., "a photo of a [CLASS]", is often directly used as the text input, and the maximum similarity between it and the input image is calculated to determine which class the image belongs to. [36, 17] apply a learnable text prompt, which consists of the class context and a learnable context and can be defined as follows:
$t_{j}=[context_{front}][CLASS_{j}][context_{end}].$ (2)
The image feature $z_{v}^{i}$ is generated by image encoder $F_{\theta_{v}}$ of input $\tilde{x}_{i}$, the text feature $z_{t}^{i}$ is generated by text encoder $F_{\theta_{t}}$ of input $t_{j}$, which can be defined as follows:
$\tilde{z}_{v}^{i}=F_{\theta_{v}}(\tilde{x}_{i}),z_{t}^{j}=F_{\theta_{t}}(t_{j}).$ (3)
@@
$\mathcal{L}(\tilde{x}_{i},t,,y_{i};F_{\theta_{v}},F_{\theta_{t}})=\sum_{j}^{N}-y_{ij}log\frac{exp(cos(\tilde{z}_{v}^{i},z_{t}^{j}))}{\sum_{j}^{N}exp(cos(\tilde{z}_{v}^{i},z_{t}^{j})},$ (4)
where the cosine similarity is applied to measure the degree of alignment of the features, and applying the softmax operation can obtain the probability that the $\tilde{x}_{i}$ aligns with the $z_{t}$.
3.2 Selecting a Longer Prompt or More Prompts in APT?
@@
From the experimental results, we are surprised to find that in the process of adversarial prompt tuning, increasing the prompt number is more effective than increasing the prompt length. Specifically, we find that using 2/4 learnable adversarial prompts of length 16 achieves better adversarial robustness compared to using a single prompt of length 32/64, with an average improvement of 3.88%/4.34%. Notably, this setup also leads to an average performance gain of 4.56%/6.43% on clean samples. Furthermore, increasing the prompt number further enhances adversarial robustness; When the prompt number increases from 2 to 4, adversarial robustness improves by an additional 0.34%. In contrast, merely increasing the prompt length yields no obvious improvement in adversarial robustness.
Figure 2: The performance of adversarial prompt tuning with different length and number on five datasets. “APT-Lm-Nk” denotes the APT with prompt length m and prompt number k. We find that increasing the prompt number can enhance more robustness than increasing the prompt length (i.e., solid lines show better performance than dotted lines).
We argue that after the total number of parameters reaches a certain level, continuing to increase the length of prompts will increase the difficulty of learning an ideal prompt, so it hardly enhances the adversarial robustness. On the contrary, shorter prompts are relatively easier to learn, and adversarial mixture prompts can generate more diverse text features, which have more possibility to align with different adversarial examples. Therefore, increasing the prompt number can further improve robustness compared with increasing prompt length.
4 Adversarial Mixture Prompt Tuning
@@ 4.1 Overall Framework
Based on the above findings, we argue that the generalization of adversarial robustness can be improved by adding adversarial mixture prompts. Therefore, we propose Adversarial Mixture Prompt Tuning (AMPT) to further improve the adversarial robustness of the VLMs. Here our optimization goal can be formulated as follows:
$\mathop{\arg\min}_{T_{m}}\mathbb{E}_{(x,t,y)\sim\mathcal{D}}(\mathcal{L}_{AMPT}(\tilde{x},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})),$ (5)
where $\mathcal{L}_{AMPT}$ denotes the optimization loss function of our AMPT, and $T_{m}$ denotes the adversarial mixture prompts, $F_{\theta_{w}}$ denotes the conditional prompt weight router. As for the adversarial examples $\tilde{x}$, we follow the “on-the-fly” setting in [17], where the attacker can access all the parameters of the VLMs including the adversarial mixture prompts but can only apply adversarial perturbations to the image $x$. And the adversarial examples $\tilde{x}$ can be formulated as follows:
$\tilde{x}=\mathop{argmax}\limits_{||\tilde{x}-x||\leq\epsilon}\mathcal{L}_{AMPT}(\tilde{x},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})),$ (6)
where $\epsilon$ denotes the maximum perturbation scale. It should be mentioned that the “on-the-fly” setting is closer to the adversarial examples in [31, 22]. For the evaluation against adversarial attacks, we also follow this type of setting as default.
@@ 4.2 Adversarial Mixture Prompts
where $t^{k}$ denotes the $k$-th learnable adversarial text prompt, which includes $N$ class text prompt: $\{t_{1}^{k},t_{2}^{k},\cdots,t_{N}^{k}\}$. Following [38, 17], the $CLASS_{j}$ context in each $t^{k}$ is represented by a sequence of class-specific vectors, and the learnable contexts are defined in the word embedding space, then $t^{k}$ for class $j$ can be formulated as follows:
$t_{j}^{k}=[V]_{1,j}^{k}...[V]_{M,j}^{k}[C_{j}],$ (8)
where the $M$ denotes the max length of learnable context. The position of $[C_{j}]$ can also be adjusted. Following [17], we apply the end position as the default position. It should be mentioned that the learnable context can also apply two different modes: unified context (UC) and class-specific context (CSC), where unified context denotes that $t_{j}^{k}$ for different classes share the same learnable context, and class-specific context denotes that $t_{j}^{k}$ for different classes has their independent learnable context.
@@
$z_{t}^{j,i}=\sum_{k}^{K}\tilde{w}_{k}^{i}F_{\theta_{t}}(t_{j}^{k}),$ (9)
where $\tilde{w}_{k}^{i}$ denotes the weights of adversarial prompt $t^{k}$ for the adversarial examples $\tilde{x}_{i}$, and $\tilde{w}_{k}^{i}$ is irrelevant to the class (not effected by $j$). $z_{t}^{j,i}$ denotes the aggregated text feature of $j$-th class for the adversarial examples $\tilde{x}_{i}$. In this way, we can obtain adversarial mixture prompts with pretty diversity through adversarial training to defend against different adversarial examples.

### v2

# MoAPT: Mixture of Adversarial Prompt Tuning for Vision-Language Models

@@ Abstract
Large pre-trained Vision Language Models (VLMs) demonstrate excellent generalization capabilities but remain highly susceptible to adversarial examples, posing potential security risks. To improve the robustness of VLMs against adversarial examples, adversarial prompt tuning methods are proposed to align the text feature with the adversarial image feature without changing model parameters. However, when facing various adversarial attacks, a single learnable text prompt has insufficient generalization to align well with all adversarial image features, which ultimately results in overfitting. To address the above challenge, in this paper, we empirically find that increasing the number of learned prompts yields greater robustness improvements than simply extending the length of a single prompt. Building on this observation, we propose an adversarial tuning method named Mixture of Adversarial Prompt Tuning (MoAPT) to enhance the generalization against various adversarial attacks for VLMs. MoAPT aims to learn mixture text prompts to obtain more robust text features. To further enhance the adaptability, we propose a conditional weight router based on the adversarial images to predict the mixture weights of multiple learned prompts, which helps obtain sample-specific mixture text features aligning with different adversarial image features. Extensive experiments across 11 datasets under different settings show that our method can achieve better adversarial robustness than state-of-the-art approaches.
@@ 1 Introduction
Large pre-trained Vision Language Models (VLMs) such as CLIP [27] have excellent generalization capabilities and can be regarded as foundation models [2] in different downstream tasks, e.g., image-text retrieval, zero-shot image classification, or image generation guidance. Due to its wide range of application scenarios, it places high requirements on security performance. However, despite its excellent performance, VLMs face many potential security risks [13, 24, 28], including the fact that visual models are vulnerable to adversarial examples [31], which can pose a serious threat to the application in actual scenarios.
@@
To eliminate this potential security risk, many works have been proposed to improve the robustness of VLMs to adversarial examples, which can be mainly divided into two types, full-parameter fine-tuning [24, 32, 29, 35] and parameter-efficient fine-tuning [24, 32, 36, 17, 39, 21]. Among them, full-parameter fine-tuning is an effective method to improve the adversarial robustness of the model. However, this method often requires a lot of computational overhead and also affects the performance of the model on general tasks. Another type of parameter-efficient method, e.g., adversarial prompt tuning [36], freezes all or most of the weights of the model and only fine-tunes some of its parameters. This type of method can also improve the adversarial robustness with lower training overhead compared with full-parameter fine-tuning, which is a promising solution. However, adversarial prompt tuning faces a serious problem: insufficient generalization. For example, for the text prompt tuning [36, 17], when only one learnable prompt is fine-tuned, the text feature is not sufficient to fit the image features for various adversarial examples, which can easily lead to overfitting and further cause potential security risks [33].
@@
Figure 1: The framework of Mixture of Adversarial Prompt Tuning (MoAPT). To enhance the adversarial robustness, we apply adversarial mixture prompt to generate diverse individual text feature, and utilize the conditional prompt weight router to obtain a sample-specific mixture text feature, and finally bring more generalization towards different adversarial examples.
@@
To enhance the generalization of the adversarial text prompt, an intuitive approach is to increase the length of the text prompt. However, we find that when it grows to a certain length, a longer prompt will bring greater optimization difficulty, and also needs higher requirements on the corresponding text encoder to deal with long prompts, finally leading to suboptimal robustness. Inspired by the Mixture of Experts (MoE) paradigm [4], we adopt an alternative strategy: increasing the number of learnable base prompts. Similar to multiple experts in MoE, we construct a composite prompt by combining several base prompts, with their weights adaptively determined based on the characteristics of the adversarial image. This approach enables the generation of more diverse and expressive text features. Moreover, since each base prompt remains short and is easier to optimize, leading to improved adversarial robustness. A preliminary experiment has empirically verified our idea that increasing the number of prompts can indeed enhance robustness more effectively than simply extending the length of the prompts.
@@
Based on the above consideration, in this paper, we propose an adversarial prompt tuning method named Mixture of Adversarial Prompt Tuning (MoAPT) to enhance the adversarial robustness of VLMs. Specifically, we fix the parameters of the text and image encoders but only optimize adversarial mixture prompts. These text prompts pass through the text encoder and generate diverse individual text features. In addition, to enhance the adaptability, we propose a conditional text weight router based on image features to predict the weights of adversarial mixture prompts and aggregate them into a sample-specific mixture text feature, so as to adaptively align with the diverse adversarial image features. A series of experiments show that our MoAPT can achieve better accuracy and robustness than state-of-the-art methods on multiple different datasets. Meanwhile, MoAPT also shows better generalization across different datasets. Our contribution can be summarized as follows:
@@
  • We find that for adversarial text prompt tuning, increasing the number of learnable text prompts can achieve a better robustness than only increasing the length of learnable text prompts within a certain range of parameters.
@@
  • We propose a novel method named Adversarial Mixture Prompt Tuning (MoAPT), which applies adversarial mixture prompts to generate diverse individual text features, where each text feature can play its unique roles for different adversarial examples, thereby alleviating the overfitting phenomenon.
@@
  • We apply a conditional text weight router based on image features to predict the weights of different text features and obtain a sample-specific mixture text feature that has pretty adaptability to align with different adversarial image features. Furthermore, we theoretically verify the effectiveness of the weight router.
@@
  • We empirically verify the effectiveness of MoAPT. Extensive experiments demonstrate that our MoAPT can outperform state-of-the-art methods against adversarial examples in adversarial robustness and generalization across different datasets.
@@ 2 Related Work
Figure 2: The performance of adversarial prompt tuning with different length and number on five datasets. “APT-Lm-Nk” denotes the APT with prompt length m and prompt number k. We find that increasing the number of prompts can enhance more robustness than increasing the prompt length (i.e., solid lines show better performance than dotted lines).
2.1 Prompt Tuning for Accuracy in VLMs
@@
2.2 Adversarial Prompt Tuning in VLMs
Due to its excellent performance and low training cost, prompt tuning has been applied to improve the robustness of VLMs. [5] applies visual prompting to enhance the adversarial robustness. Furthermore, TeCoA [24] and PMG-AFT [32] employ visual prompt tuning to improve the adversarial robustness of VLMs. AdvPT [36] and APT [17] are proposed to apply text prompt tuning to further enhance the VLMs against image attacks. FAP [39] tries to enhance the robustness via bimodal tuning, while APD [21] further extends FAP into the adversarial distillation setting. To solve the insufficient generalization, [33] applies Test-Time Adversarial Prompt Tuning (TAPT) to learns defensive bimodal (textual and visual) prompts during testing process. Different from the above research, this paper improves adversarial robustness through adversarial mixture prompt tuning during training process, which tries to solve the existing issue from another view but not conflict with each other.
3 The Necessity of Mixture Prompts
3.1 Formulation of Adversarial Prompt Tuning
@@
where $x$ and $t$ are the image and text pairs belong to the dataset $\mathcal{D}$. For the image classification task with $N$ classes, texts $t$ also contain $N$ different prompts: $\{t_{1},t_{2},\cdots,t_{N}\}$. $\tilde{x}$ denotes adversarial examples. $y$ denotes the ground truth. $y_{in}$ indicates whether the image $x_{i}$ and text $t_{n}$ pair match, if the image $x_{i}$ and text $t_{n}$ match, $y_{in}$ is equal to 1, otherwise $y_{in}$ is equal to 0; the $F_{\theta_{v}}$ and $F_{\theta_{t}}$ are the image encoder and text encoder of CLIP.
Meanwhile, as for the text $t$, a fixed text template, e.g., "a photo of a [CLASS]", is often directly used as the text input, and the maximum similarity between it and the input image is calculated to determine which class the image belongs to. [36, 17] apply a learnable text prompt, which consists of the class context and a learnable context as follows:
$t_{n}=[context_{front}][CLASS_{n}][context_{end}].$ (2)
The image feature $z_{v}^{i}$ is generated by image encoder $F_{\theta_{v}}$ of input $\tilde{x}_{i}$, the text feature $z_{t}^{n}$ is generated by text encoder $F_{\theta_{t}}$ of input $t_{n}$, which can be defined as follows:
$\tilde{z}_{v}^{i}=F_{\theta_{v}}(\tilde{x}_{i}),z_{t}^{n}=F_{\theta_{t}}(t_{n}).$ (3)
@@
$\mathcal{L}(\tilde{x}_{i},t,y_{i};F_{\theta_{v}},F_{\theta_{t}})=-\sum_{n=1}^{N}y_{in}log\frac{exp(cos(\tilde{z}_{v}^{i},z_{t}^{n}))}{\sum_{m=1}^{N}exp(cos(\tilde{z}_{v}^{i},z_{t}^{m})},$ (4)
where the cos similarity is applied to measure the alignment degree of the features, and applying the softmax operation can obtain the probability that the $\tilde{x}_{i}$ aligns with the $z_{t}$.
3.2 A Longer Prompt or More Prompts in APT?
@@
The results surprisingly reveal that, during adversarial prompt tuning, increasing the number of prompts is more effective than increasing their length. Specifically, we find that using 2/4 learnable adversarial prompts of length 16 achieves better adversarial robustness compared to using a single prompt of length 32/64, with an average improvement of 3.88%/4.34%. Notably, this setup also leads to an average performance gain of 4.56%/6.43% on clean samples. Furthermore, increasing the number of prompts further enhances adversarial robustness; When the number of prompts increases from 2 to 4, adversarial robustness improves by an additional 0.34%. In contrast, merely increasing the prompt length yields no obvious robustness improvement.
We argue that after the total number of parameters reaches a certain level, continuing to increase the length of prompts will increase the difficulty of learning an ideal prompt. On the contrary, shorter prompts are relatively easier to learn, and adversarial mixture prompts can generate more diverse text features, which have more possibility to align with adversarial examples. Therefore, increasing the number of prompts can further improve robustness compared with increasing prompt length.
4 Mixture of Adversarial Prompt Tuning
@@ 4.1 Overall Framework
Based on the above findings, we argue that the generalization of adversarial robustness can be improved by adding adversarial mixture prompts. Therefore, we propose Mixture of Adversarial Prompt Tuning (MoAPT) to further improve the adversarial robustness of the VLMs. Here our optimization goal can be formulated as follows:
$\mathop{\arg\min}_{T_{m},\theta_{w}}\mathbb{E}_{(x,t,y)\sim\mathcal{D}}(\mathcal{L}_{MoAPT}(\tilde{x},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})),$ (5)
where $\mathcal{L}_{MoAPT}$ denotes the optimization loss function of our MoAPT, and $T_{m}$ denotes the adversarial mixture prompts, $F_{\theta_{w}}$ denotes the conditional prompt weight router. As for the adversarial examples $\tilde{x}$, we follow the “on-the-fly” setting in [17], where the attacker can access all the parameters of the VLMs including the adversarial mixture prompts but can only apply adversarial perturbations to the image $x$. And the adversarial examples $\tilde{x}$ can be formulated as follows:
[…]
where $\epsilon$ denotes the maximum perturbation scale. It should be mentioned that the “on-the-fly” setting is closer to the adversarial examples in [31, 22], which can access all parameters of the model and only modify the images. For the evaluation against adversarial attacks, we also follow this type of setting.
@@ 4.2 Adversarial Mixture Prompts
where $t^{k}$ denotes the $k$-th learnable adversarial text prompt, which includes $N$ class text prompt: $\{t_{1}^{k},t_{2}^{k},\cdots,t_{N}^{k}\}$. Following [38, 17], the $CLASS_{n}$ context in each $t^{k}$ is represented by a sequence of class-specific vectors, and the learnable contexts are defined in the word embedding space, then $t^{k}$ for class $n$ can be formulated as follows:
$t_{n}^{k}=[V]_{1,n}^{k}...[V]_{M,n}^{k}[C_{n}],$ (8)
where the $M$ denotes the max length of learnable context. The position of $C_{n}]$ can also be adjusted. Following [17], we apply the end position as the default position.
@@
$z_{t}^{n,i}=\sum_{k=1}^{K}\tilde{w}_{k}^{i}F_{\theta_{t}}(t_{n}^{k}),$ (9)
where $\tilde{w}_{k}^{i}$ denotes the weights of adversarial prompt $t^{k}$ for the adversarial examples $\tilde{x}_{i}$, and $\tilde{w}_{k}^{i}$ is irrelevant to the class (not effected by $n$). $z_{t}^{n,i}$ denotes the mixture text feature of $n$-th class for the adversarial examples $\tilde{x}_{i}$. In this way, we can obtain adversarial mixture prompts with pretty diversity through adversarial training to defend against different adversarial examples.